OpenStudy (anonymous):

The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76. If x represents the width of the rectangle, then which expression represents the perimeter? i got 2x + 2?

myininaya (myininaya):

length is 2 more than twice the width L = 2 + 2* x so one equation is L=2+2x the perimeter of a rectangle is found by adding up all the sides so you have x+x+L+L=76 2x+2L=76 ------ if you want to find the dimensions (use the equations i have above) 2x+2(2+2x)=76 2x+4+4x=76 6x+4=76 6x=76-4 6x=72 x=12 is the width so the length is L=2+2x=2+2(12)=2+24=26